Airplane normally fly at altitude of 36000 ft. Find the atmospheric pressure at that altitude.


1
Expert's answer
2022-04-01T16:32:52-0400

The standard temperature T=288.15K

 corresponds to the International Standard Atmosphere (ISA) model.

"P=P_0exp(-\dfrac{Mg}{RT}h)"

Average sea level pressure=101,325 kPa "(P_0)"

Gravitational acceleration= 9.807 m/s2 (g)

Molar mass of air=0.02896 (kg/mol) (M)

The standard temperature T=288.15K

Universal Gas Constant=8.3143 (N*m)/(mol*K) (R)

36000ft=10970m

101,325*exp(-0.00012*10970) kPa=27.1651kPa